# EEG Emotion Copilot: Pruning LLMs for Emotional EEG Interpretation with Assisted Medical Record Generation

### v1

@@ Abstract
In the fields of affective computing (AC) and brain-machine interface (BMI), the analysis of physiological and behavioral signals to discern individual emotional states has emerged as a critical research frontier. While deep learning-based approaches have made notable strides in EEG emotion recognition, particularly in feature extraction and pattern recognition, significant challenges persist in achieving end-to-end emotion computation, including real-time processing, individual adaptation, and seamless user interaction. This paper presents the EEG Emotion Copilot, a system leveraging a lightweight large language model (LLM) operating in a local setting. The system is designed to first recognize emotional states directly from EEG signals, subsequently generate personalized diagnostic and treatment suggestions, and finally support the automation of electronic medical records. The proposed solution emphasizes both the accuracy of emotion recognition and an enhanced user experience, facilitated by an intuitive interface for participant interaction. We further discuss the construction of the data framework, model pruning, training, and deployment strategies aimed at improving real-time performance and computational efficiency. Privacy concerns are also addressed, with a focus on ethical data collection, processing, and the protection of users’ personal information. Through these efforts, we aim to advance the application of AC in the medical domain, offering innovative approaches to mental health diagnostics and treatment.
@@
Lightweight LLM, Model pruning, Emotion recognition, Assisted electronic medical record
@@ I Introduction
The application of affective computing (AC) in brain-machine interfaces (BMI) is emerging as a key research frontier. Affective computing [1] seeks to identify emotional states by analyzing physiological and behavioral signals, such as electroencephalograms (EEG) [2] [3], heart rate [4], facial expressions [5], and voice [6]. Advances in this field have expanded the possibilities for BMI technology, particularly in enhancing human-computer interaction [7], and increasing its potential for practical applications in medical rehabilitation [8] and other areas [9].
The powerful learning capabilities of deep neural networks (DNNs) have driven rapid advancements in AC for emotion recognition via EEG signals [10], with numerous deep learning approaches proposed [11], [12], [13]. However, achieving end-to-end EEG-based AC remains challenging. Large language models (LLMs) [14], [15] built on the Transformer architecture [16] have demonstrated strong contextual understanding [17]. Models such as EEG-GPT [18] and ChatGPT-BCI [19] focus primarily on brain state identification but have yet to effectively address diverse, human-centered tasks. The integration of LLMs with EEG signals to generate personalized diagnoses, treatment plans, and assist in electronic medical records remains an urgent and complex challenge with the following limitations.
Firstly, building a suitable human-based corpus from diverse databases for LLMs is a challenging task. While public datasets like Wikitext [20] and C4 [21], along with specialized datasets like SEED [22] and FACED [23], offer substantial prior knowledge, this information remains fragmented. The challenge lies in linking these datasets to create a unified data structure that facilitates language model development and enables accurate responses through appropriate prompts.
Secondly, EEG signals can be viewed as a form of long text data [24], yet the signal sequences from a single channel often exhibit significant redundancy for current [25], [26] and traditional machine learning approaches [27], [28]. LLMs require a considerable number of tokens [16] to process this redundancy, resulting in unnecessary computational overhead. Additionally, many AC tasks rely on data from 32, 64, or more channels, further diminishing computational efficiency and presenting challenges for real-time performance. Therefore, effective data compression is critical to optimizing lightweight LLMs for real-time applications.
@@
Therefore, we propose EEG Emotion Copilot, an intelligent system based on a lightweight, locally-running language model that utilizes EEG signals to perform emotion status recognition and subsequently generate corresponding diagnosis and treatment plans. This aims to automatically create assisted electronic medical records, as illustrated in Figure 1. In the following sections, we will detail the construction of an effective data framework, the process of model pruning, model training, and final model deployment.
Figure 1: Flowchart of EEG Emotion Copilot. The original EEG signal is preprocessed and transformed via wavelet to shorten the signal length. The final prompt is constructed using the initialized prompt, demographic data, emotional label, and treatment as training data. Using the Qwen2-0.5B-pretrained model [29] as an example, we prune it by 50% and restore its performance with Wikitext [20], C4 [21], and other datasets to obtain the pruned Qwen2-0.5B model. During the Model Fine-tuning stage, a Warm-up phase gradually increases the learning rate, and Lora [30] is used for fine-tuning. Finally, the model is deployed, utilizing the RAG (Retriever-Reader-Generator) technique [31] to enhance retrieval performance, and a dialogue method in the graphical interface to improve interactivity.
@@ II RELATED WORK
II-A Data Structure in Prompting LLMs
@@
For specific tasks in LLM prompt engineering, tailored prompts are essential to obtain satisfactory answers [32]. Although datasets like Wikitext [20] and C4 [21] contain extensive knowledge, constructing specific datasets is necessary for optimal performance in particular tasks. For participation-centered tasks, incorporating more personal information can yield better answers. In much of the work on AC, researchers focus on distinguishing emotions using convolutional neural networks (CNNs) or recurrent neural networks (RNNs) [33], [34], with EEG signals primarily used to assess model quality [35], [36], [37], [38], [39]. The excessive length of input EEG data is often overlooked, leading to low memory and computational costs. However, in practical deployment scenarios, the EEG data length significantly impacts reasoning effectiveness [40], [41]. LLMs must split text sequences into tokens for processing, with computational and memory requirements increasing exponentially with input length. This poses challenges for LLMs deployment, especially on low-resource devices or in real-time scenarios. Although data chunking strategies [42] can split long sequences for training and deployment, our experiments show limited effectiveness in practice. Segmentation increases data volume and complexity, complicating category identification. Therefore, a reasonable data compression process is necessary to LLMs.
II-B Model Pruning in LLMs
@@
Model pruning [43] has been widely applied in DNNs to reduce computational complexity and mitigate model hallucinations [44]. Frankle and Carbin’s lottery ticket hypothesis [45] posits that small subnetworks, when trained from scratch, can achieve comparable performance. Li et al. [46] and Han et al. [47] introduced filter and weight pruning techniques, while Michel et al. [48] explored attention head pruning in Transformer models. Zhuang et al. [49] employed gradient-based structured pruning, and Liu et al. [50] proposed MetaPrune, which combines pruning with neural architecture search (NAS). Guo et al. [51] focused on channel pruning for efficient inference. Comprehensive surveys by Blalock et al. [52] and He et al. [53] have highlighted the effectiveness of various pruning techniques and outlined future research directions. In the context of LLMs, methods such as SparseGPT [54] and LLM-Pruner [55], primarily based on the LLaMA model [56], have demonstrated promising results on public datasets and claim faster performance on local machines. However, the hardware used for testing, such as RTX 4090 GPUs, significantly outperforms standard PCs, leaving the inference cost still relatively high. Thus, for local computations with limited resources, there remains a need for more efficient and feasible pruning methods to further reduce computational costs.
II-C Assisted Electronic Medical Record Generation
@@
Electronic medical records (EMRs) date back to the 1960s [57]. With advancements in computer science and artificial intelligence (AI), EMRs have evolved from merely storing and managing patient information to incorporating more complex functionalities. Today, AI models are increasingly integrated into diagnosing mental illnesses and detecting internal organ pathologies, enhancing the capabilities of EMRs. These models can be embedded within ”assisted electronic medical record” shown in Figure 2, where multiple copilots collaborate to process diagnostic data and produce accurate results. For simpler tasks, a single copilot may suffice. In response, we developed a new EEG Emotion Copilot that leverages a lightweight local LLM to perform emotion recognition based on EEG signals and provide corresponding diagnosis and treatment plans.
@@
Figure 2: Schematic diagram of the assisted electronic medical record generated by intelligent copilots. It includes Basic information, Medical History, Physical Examination Records, Diagnosis Information, Treatment Plan, Laboratory Results, Follow-up Records, and Medical Expenses. Multiple copilots diagnose symptoms in corresponding fields and give comprehensive diagnosis results combined with laboratory results, and provide treatment plans.
@@ III-A Data Structures of Prompt in EEG Emotion Copilot
Fine-tuned LLMs often achieve strong performance and generate end-to-end outputs, particularly for simple classification tasks. For example, distinguishing whether a subject’s emotional state is positive or negative based on EEG data is relatively straightforward. However, distinguishing more nuanced emotions, such as those in the nine-class FACED dataset [23], or accurately identifying emotions in real-world scenarios, remains challenging. To address this, we incorporate demographic features, including gender, age, and facial characteristics, into the data. General LLMs are typically trained on broad datasets like Wikitext [20] and Common Crawl [21] for general knowledge and logical reasoning. For domain-specific applications, they are trained on specialized datasets, such as BioBERT [58] for biomedicine, ClinicalBERT [59] for medical texts, and FinBERT [60] for financial texts. These datasets often consist of declarative sentences, but for tasks like logical reasoning, data are stored as question-answer pairs, as shown in Figure 3. For EEG-based emotion diagnosis and treatment planning, a question-answer system offers a more effective solution.
Figure 3: Comparison of three data structures: the general data structure (left), the question-answering system data structure (middle), and the data structure used in the proposed EEG Emotion Copilot (right).
In the data structure of the proposed EEG Emotion Copilot, we initialize the system with the prompt showed in the right part of Figure 2: ”You are an EEG emotion analyzer. I will input the patient’s personal information and EEG signals from specific electrode positions. Please infer the patient’s current emotional state and provide a detailed diagnosis along with personalized treatments based on your knowledge base.” Following initialization, we input the subject’s demographic data, including gender, age, and facial features, to assess their emotional state. Next, the acquired EEG signals are processed. As illustrated in Figure 1, preprocessing steps such as artifact removal and re-referencing are applied to the signals, followed by wavelet transformation for signal compression. These processed signals are then used as input. Finally, the structure is completed by adding emotion labels and corresponding treatment plans in the second and third sections, respectively.
@@ III-B Model Pruning in EEG Emotion Copilot
Although some pruning schemes may achieve better results, we use torch pruning [61] in this paper for its flexibility and versatility. We set the pruning_ratio to 0.5, reducing the model’s parameters by half. To restore performance, we train and fine-tune the pruned model on a public dataset. We then pseudo-quantize the model weights by pseudo-quantizing [62] the top 1% of the most important weights, identified by their calculated “importance.” The remaining weights are left unchanged. For details, please see Figure 4.
Figure 4: Flowchart of model pruning in EEG Emotion Copilot: The upper left presents a rough framework diagram, while the right side illustrates torch pruning with a 0.5 pruning ratio. In this process, the embed_tokens in the embedding layer is reduced by half, and the input_feature and out_feature of Qwen2SdpaAttention and Qwen2MLP are halved compared to the original model. The final lm_head output remains unchanged, as it is directly tied to the vocabulary size. The lower left depicts a simplified torch pruning process.
Figure 5: Flowchart of electronic medical record generation implemented in EEG Emotion Copilot.
@@ III-C Model Fine-tuning in EEG Emotion Copilot
Even when a pre-trained model is fine-tuned on a public dataset post-pruning, certain tasks may still require fine-tuning with a proprietary dataset. In this study, the Low-Rank Adaptation (LoRA) [30] has been employed, which introduces low-rank matrices in specific layers to adjust parameters, avoiding the need for complete retraining of all weights, thus enabling faster and more efficient adaptability. Additionally, a warm-up strategy is often applied at the onset of training to facilitate convergence and improve overall model performance.

### v2

# EEG Emotion Copilot: Optimizing Lightweight LLMs for Emotional EEG Interpretation with Assisted Medical Record Generation

@@ Abstract
In the fields of affective computing (AC) and brain-machine interface (BMI), the analysis of physiological and behavioral signals to discern individual emotional states has emerged as a critical research frontier. While deep learning-based approaches have made notable strides in EEG emotion recognition, particularly in feature extraction and pattern recognition, significant challenges persist in achieving end-to-end emotion computation, including real-time processing, individual adaptation, and seamless user interaction. This paper presents the EEG Emotion Copilot, a system optimizing a lightweight large language model (LLM) with 0.5B parameters operating in a local setting, which first recognizes emotional states directly from EEG signals, subsequently generates personalized diagnostic and treatment suggestions, and finally supports the automation of assisted electronic medical records. Specifically, we demonstrate the critical techniques in the novel data structure of prompt, model pruning and fine-tuning training, and deployment strategies aiming at improving real-time performance and computational efficiency. Extensive experiments show that our optimized lightweight LLM-based copilot achieves an enhanced intuitive interface for participant interaction, superior accuracy of emotion recognition and assisted electronic medical records generation, in comparison to such models with similar scale parameters or large-scale parameters such as 1.5B, 1.8B, 3B and 7B. In summary, through these efforts, the proposed copilot is expected to advance the application of AC in the medical domain, offering innovative solution to mental health monitoring. The codes will be released at https://github.com/NZWANG/EEG_Emotion_Copilot.
@@
Lightweight LLM, Model pruning, Model fine-tuning, Emotion recognition, Assisted electronic medical record
@@ I Introduction
THE application of affective computing (AC) in brain-machine interfaces (BMI) is emerging as a key research frontier. Affective computing [1] seeks to identify emotional states by analyzing physiological and behavioral signals, such as electroencephalograms (EEG) [2] [3], heart rate [4], facial expressions [5], voice [6], etc. Advances in this field have expanded the possibilities for BMI technology, particularly in enhancing human-machine interaction [7], [8], and increasing its potential for practical applications in rehabilitation [9], [10] and other areas [11].
The powerful capabilities of deep neural networks (DNNs) have driven rapid advancements in AC for emotion recognition via EEG signals [12], [13], with numerous deep learning approaches proposed [14], [15]. Recently, large language models (LLMs) [16], [17] built on Transformer architecture [18] have demonstrated strong contextual understanding [19]. Models such as EEG-GPT [20] and ChatGPT-BCI [21] focus primarily on brain state identification but have yet to effectively address diverse, human-centered tasks. The integration of LLMs with EEG signals to generate personalized diagnoses, treatment plans, and assisted electronic medical records remains an urgent and complex challenge with the following limitations.
Figure 1: Flowchart of EEG Emotion Copilot. The original EEG signal is preprocessed and transformed via wavelet to shorten the signal length. The final prompt is constructed using the initialized prompt, demographic data, emotional label, and treatment as training data. Using the Qwen2-0.5B-pretrained model [22] as an example, we prune it with the radio of 50% and fine-tuning the pruned model directly with specific dataset. During the Model Fine-tuning stage, a warm-up phase gradually increases the learning rate, and LoRA [23] is used for fine-tuning. Finally, the model is deployed, utilizing the RAG (Retriever-Reader-Generator) technique [24] to enhance retrieval performance, and a dialogue deployment of the graphical interface to improve interactivity.
Firstly, building a suitable human-based multimodal corpus from diverse databases for LLMs is a challenging task. While public datasets like Wikitext [25] and Common Crawl [26], along with specialized datasets like SEED [27] and FACED [28], offer substantial prior knowledge, this information remains fragmented. The challenge lies in linking these datasets to create a unified data structure that facilitates language model development and enables accurate responses through appropriate prompts.
Secondly, EEG signals can be viewed as a form of long text data [29], yet the signal sequences from a single channel often exhibit significant redundancy for machine learning approaches [30], [31] [32], [33]. LLMs require a considerable number of tokens [18] to process this redundancy, resulting in unnecessary computational overhead. Additionally, many AC tasks rely on EEG data from 32, 64, or more channels, further diminishing computational efficiency and presenting challenges for real-time performance. Therefore, effective data compression is critical to optimizing lightweight LLMs for real-time applications.
@@
Therefore, we propose EEG Emotion Copilot, an intelligent system based on a lightweight, locally-running language model that utilizes EEG signals to first perform emotion recognition, subsequently generate corresponding diagnosis and treatment plans, and finally create assisted electronic medical records, as illustrated in Fig. 1.
@@ II RELATED WORK
II-A Lengthy EEG Signal Embedded in Prompt
@@
For specific tasks in LLM prompt engineering, tailored prompts are essential to obtain reasonable answers [34]. Although datasets like Wikitext [25] and Common Crawl [26] contain extensive knowledge, constructing specific datasets is necessary for optimal performance in particular tasks. In most studies on AC, researchers focus on distinguishing emotions using convolutional neural networks (CNNs) or recurrent neural networks (RNNs) [35], with EEG signals primarily used to assess model quality [36], [37]. Recently, LLMs should split EEG sequences into tokens for processing, with computational and memory requirements increasing exponentially with input length [38]. This poses challenges for LLMs deployment, especially on low-resource devices or in real-time scenarios. Although data chunking strategies [39] can split long sequences for training and deployment, the EEG data length significantly impacts reasoning effectiveness [40], [41]. Therefore, a reasonable data compression method is necessary to LLMs.
II-B Model Pruning in Deep Learning
@@
The Frankle and Carbin’s lottery ticket hypothesis [42] posits that small subnetworks, when trained from scratch, can achieve comparable performance. Based on it, model pruning [43] has been widely applied in DNNs to reduce computational complexity and mitigate model hallucinations [44]. For instance, Li et al. [45] and Han et al. [46] introduced filter and weight pruning techniques, while Michel et al. [47] explored attention head pruning in Transformer models. Zhuang et al. [48] employed gradient-based structured pruning, and Liu et al. [49] proposed MetaPrune, which combines pruning with neural architecture search (NAS). Guo et al. [50] focused on channel pruning for efficient inference. Comprehensive surveys by Blalock et al. [51] and He et al. [52] have highlighted the effectiveness of various pruning techniques and outlined future research directions. In the context of LLMs, methods such as SparseGPT [53] and LLM-Pruner [54], primarily based on the LLaMA model [55], have demonstrated promising results on public datasets and claim faster performance on local machines. However, the hardware used for inference, such as RTX 4090 GPUs, significantly outperforms standard PCs, leaving the inference cost still relatively high. Thus, for local computations with limited resources, there remains a need for more efficient and feasible pruning methods to further reduce computational costs.
II-C Assisted Electronic Medical Records
@@
Electronic medical records (EMRs) date back to the 1960s [56]. With advancements in computer science and artificial intelligence (AI), EMRs have evolved from merely storing and managing patient information to incorporating more complex functionalities. Currently AI models are increasingly integrated into diagnosing mental illnesses and detecting internal organ pathologies, enhancing the capabilities of assisted EMRs. The sample of ”assisted electronic medical record” is shown in Fig. 2, where the intelligent copilot can process the multimodal diagnostic data and produce accurate results. In this study, we developed a new EEG Emotion Copilot that leverages a lightweight local LLM to perform emotion recognition based on EEG signals, and to further generate the corresponding assisted EMR.
@@
Figure 2: Schematic diagram of the assisted electronic medical record generated by intelligent copilot. It can include Basic Information, Medical History, Physical Examination Records, Diagnosis Information, Treatment Plan, Laboratory Results, Follow-up Records, and Medical Expenses.
@@ III-A Data Structures of Prompt in EEG Emotion Copilot
Fine-tuned LLMs often achieve strong performance to generate end-to-end outputs, particularly for simple classification tasks. For instance, distinguishing whether a subject’s emotional state is positive or negative based on EEG data is relatively straightforward. However, distinguishing more nuanced emotions, such as nine-class emotions in FACED dataset [28], or accurately identifying emotions in real-world scenarios, remains challenging.
For Wikitext or similar datasets, these datasets often consist of declarative sentences (left part of Fig. 3), but for tasks like logical reasoning, data are stored as question-answer pairs, as shown in the middle part of Fig. 3. Despite these advancements, current data structures encounter significant limitations when applied to EEG-based emotion diagnosis and treatment planning. These challenges arise particularly in integrating multimodal data and addressing task-specific requirements. To address these gaps, we propose a novel structure leveraging a question-answer system tailored to EEG-related tasks.
Figure 3: Comparison of three data structures as to prompt in LLMs: the general data structure (left), the question-answering system data structure (middle), and the data structure used in the proposed EEG Emotion Copilot (right).
In the data structure of the proposed EEG Emotion Copilot, we initialize the system with the prompt showed in the right part of Fig. 3: ”You are an EEG emotion analyzer. I will input the patient’s personal information and EEG signals from specific electrode positions. Please infer the patient’s current emotional state and provide a detailed diagnosis along with personalized treatments based on your knowledge base.” Following initialization, we input the subject’s demographic data, including gender, age, and facial features, to assess their emotional state. Next, the acquired EEG signals are processed. As illustrated in Fig. 1, preprocessing steps such as artifact removal and re-referencing are applied to the signals, followed by wavelet transformation for signal compression. These processed signals are then used as input. Finally, the structure is completed by adding emotion labels and corresponding treatment plans in the second and third sections, respectively.
@@ III-B Model Pruning in EEG Emotion Copilot
In this study, for the flexibility and versatility, we applied the torch pruning [57] to perform the model pruning task with the pruning_ratio equal to 0.5, and the detailed process is showed in Fig. 4.
Firstly, based on the architecture of the Qwen2-0.5B model, the total number of layers $L$ is determined as follows:
$\mathcal{L}=Num(Qwen2\_layer)=Num(Embedding\_layer)$ (1)
$+Num(Qwen2Decoder\_layer)+Num(Lm\_head)=25.$
The input and output of each layer are represented as $f_{i}^{-}$ and $f_{j}^{+}$, respectively, where $0\leq i\leq 25,0\leq j\leq 25$ and $i,j\in\mathbb{N}^{+}$.
Next, we construct a layer dependency graph $\mathcal{F}$ according to [57]. Based on the symmetric mechanism of input $f_{i}^{-}$ and output $f_{j}^{+}$, the dependency graph $\mathcal{F}$ is defined as a $\mathcal{L}\times 2\mathcal{L}$ symmetric matrix, specifically as follows:
$\mathcal{F}\left(f_{i}^{-},f_{j}^{+}\right)$ $=\mathcal{F}\left(f_{j}^{+},f_{i}^{-}\right)$ (2)
$=$ $\begin{cases}1,&\text{if }f_{i}^{-}\leftrightarrow f_{j}^{+}\text{ or }\\ &\;\left(i=j\text{ and }\operatorname{sch}\left(f_{i}^{-}\right)=\operatorname{sch}\left(f_{j}^{+}\right)\right),\\ 0,&\text{otherwise}\end{cases}$
where $\text{sch}(p)$ operation is used to determine which parameters in the parameter group $p$ need to be retained. Formula 2 returns a boolean value, where $\text{sch}(f_{i}^{+})=\text{sch}(f_{j}^{-})$ indicates that the two parameter groups are subject to the same pruning scheme.
In addition, we use the $\mathcal{L}_{2}$ norm to calculate the importance of the parameter group corresponding to each layer [57]. Based on the Qwen2 architecture, we calculate the importance $\mathcal{I}$ of the Embedding layer, Self-Attention layer, MLP layer, and Linear layer in turn as follows:
$\small\mathcal{I}(\text{embed\_tokens})=\sum_{w\in\text{embed\_tokens}}\|w\|_{2}^{2},$ (3)
$\footnotesize\mathcal{I}(s_{i})=\sum_{w\in s_{i}}\|w\|_{2}^{2},\;S=\{s_{i}\}=\{q_{\text{proj}},k_{\text{proj}},v_{\text{proj}},o_{\text{proj}}\},$ (4)
$\footnotesize\mathcal{I}(m_{i})=\sum_{w\in m_{i}}\|w\|_{2}^{2},\;M=\{m_{i}\}=\{gate_{\text{proj}},up_{\text{proj}},down_{\text{proj}}\},$ (5)
$\small\mathcal{I}(\text{lm\_head})=\sum_{w\in\text{lm\_head}}\|w\|_{2}^{2}.$ (6)
Based on the above formula, we first group the parameters $\mathcal{P}$ to get set $\mathcal{P^{\prime}}$, as shown in Algorithm 1. Then, we proceed to prune the model, as shown in Algorithm 2.
Figure 4: Flowchart of model pruning in EEG Emotion Copilot: The upper left presents a rough framework diagram, while the right side illustrates torch pruning with a 0.5 pruning ratio. In this process, the input_feature and out_feature of Qwen2SdpaAttention and Qwen2MLP are halved compared to the original model. The final output of lm_head remains unchanged, as it is directly tied to the vocabulary size. The lower left depicts a simplified torch pruning process.
Algorithm 1 Parameter Grouping
1:Input: Dependency graph $\mathcal{F}$, parameter group set $\mathcal{P}$
2:Output: Group set $\mathcal{P}^{\prime}$
3:Initialize $\mathcal{P}^{\prime}\leftarrow\emptyset$
4:for all $i$in $\mathcal{L}$  do
5:     Create new group $p\leftarrow\{k\}$
6:     Expand group $p$:
7:     while there exists an unvisited node $j$ such that $\mathcal{F}$$(f_{i}^{-},f_{j}^{+})=1$ do
8:         $p\leftarrow p\cup\{j\}$
9:         Mark $j$ as visited
10:     end while
11:     $\mathcal{P}^{\prime}\leftarrow\mathcal{P}^{\prime}\cup p$
12:end for
13:return $\mathcal{P}^{\prime}$
@@ III-C Model Fine-tuning in EEG Emotion Copilot
To restore the performance of the pruned model $\mathcal{M^{\prime}}$, we explored two fine-tuning strategies after applying torch-pruning [57], as illustrated in Fig. 5. Strategy 1 is only training the pruned model on the specific EEG dataset to focus the model on task-relevant features with Low-Rank Adaptation (LoRA) [23]. Strategy 2 firstly involves full fine-tuning the pruned model on public datasets Wikitext [25] and Common Crawl [26], to leverage their diverse categories to establish generalized features, and subsequently fine-tuning on the specific EEG dataset with LoRA. Moreover, a warm-up strategy is applied at the beginning of training to facilitate convergence and improve overall model performance.
Specifically, the model’s fine-tuning process on the specific EEG dataset is illustrated in Fig. 6, dividing the training process into five steps, each with three epochs. In steps 1–4, we used 250 Hz data from 55 subjects of FACED [28]. Interestingly, the model’s performance in step 3 surpassed that in step 4, despite minimal differences in training_loss, suggesting that excessive training may be unnecessary and even detrimental for lightweight models. In step 5, 1000 Hz data from 68 FACED subjects pre-trained on step 4 yielded richer responses, confirming the model’s capability to learn the data structure and generate task-specific answers effectively.